OpenStudy (anonymous):

what is the curvature of y=x"3

OpenStudy (amistre64):

k=1/r sounds familiar

OpenStudy (amistre64):

all thats needed now is to decipher that notation ....

OpenStudy (anonymous):

give me more detail about it plz???

OpenStudy (amistre64):

the curvature of a function at a point relates to the radius of the osculating circle (tangent circle?) at a given point.

OpenStudy (amistre64):

without knowing anything else about the problem posted, its hard to be any more descriptive

OpenStudy (anonymous):

sir kindly solve it

OpenStudy (amistre64):

i would if i could figure out what y=x"3 meant.

OpenStudy (anonymous):

sir its mean y=x(power 3)

OpenStudy (anonymous):

Find the curvature of y = x 3???

OpenStudy (amistre64):

ahh, x^3

OpenStudy (amistre64):

\[k=\frac{|y''|}{(1+(y')^2~)^{3/2}}\]

OpenStudy (amistre64):

y = x^3 y' = 3x^2 y'' = 6x \[k=\frac{|6x|}{(1+9x^4)^{3/2}}\]

OpenStudy (anonymous):

thanx thanx sir (y)

OpenStudy (amistre64):

youre welcome